OpenStudy (anonymous):

find the vector v with the given magnitude and the same direction ||v||=10, u=2i-3j

OpenStudy (shubhamsrg):

Hint: vector is parallel to u, that means your vector is of the form ku, where k is any constant.

OpenStudy (dumbcow):

Find unit vector of "u" then multiply vector by ||v|| \[||u|| = \sqrt{2^{2}+(-3)^{2}} = \sqrt{13}\] unit vector: \[\frac{2}{\sqrt{13}}i - \frac{3}{\sqrt{13}}j\] multiply by 10: \[v = \frac{20}{\sqrt{13}}i - \frac{30}{\sqrt{13}}j\] rationalize denominators \[v = \frac{20\sqrt{13}}{13}i - \frac{30\sqrt{13}}{13}j\]